# Asymmetric Errors

### v2

@@ 1.1 Background and motivation
Results in particle physics are often given with errors which are asymmetric, of the form ${x}^{+{\sigma^{+}}}_{-{\sigma^{-}}}$. For instance, taking some results from the latest (at the time of writing) EPS conference:
@@
    from maximum likelihood estimation with errors given by the $\Delta\ln L=-{1\over 2}$ method, and the log likelihood is not a symmetric parabola. This is also shown in Figure 1. (The $\Delta\ln L=-{1\over 2}$ method is an approximation, but its use is widespread, essentially universal, and in what follows this is what is meant by ’likelihood errors’.)
@@
The third case does not hold conceptual problems as one has, in principle, full knowledge of the underlying distribution function. The challenges come with the first and second cases, where one has to handle the quoted value and errors with no further information about how they arose. This leads to us consider more closely some fundamental facts about the nature of ‘errors’ which are normally hidden from view by the convenient properties of the Gaussian function.
@@ 1.2 Beyond the Gaussian
The difference between rms spread ‘pdf errors’ and confidence interval ‘likelihood errors’ requires clarification. They both refer to uncertainties, the former from random probability and the latter from statistical inference. The Poisson process provides a simple illustration (though for a discrete probability distribution rather than a continuous pdf): the former describes the uncertainty of predicting results from a particular true mean, the latter the uncertainty of inferring the true value from a particular result.
@@
Pdf errors arise especially through measurements and their errors in experimental work. At high school and university the physicist is firmly taught never to quote an experimental result as a single number $x$, but always to give its error $\sigma$. That $\sigma$ is the standard deviation of the Gaussian distribution, as given in (1), from which this value is a sample. If, for better accuracy, the mean of $N$ measurements is quoted then it is accompanied by the standard error $\sigma/\sqrt{N}$ which is the standard deviation of the pdf of $\overline{x}$. The well-known formula for the combination of errors, given below as Equation 2, known to statisticans as the delta method, builds on this. It describes the combination of measurement errors: if $I$ is sampled from a Gaussian with standard deviation $\sigma_{I}$, and $V$ with $\sigma_{V}$, then the resistance measurement $R=V/I$ is sampled from a Gaussian for which (assuming the function can legitimately be linearised) $\sigma_{R}^{2}=\sigma_{V}^{2}/I^{2}+V^{2}\sigma_{I}^{2}/I^{4}$. The quoted result, $V/I\pm\sigma_{R}$, is saying the value has been drawn from a Gaussian distribution of unknown mean and standard deviation $\sigma_{R}$.
Such measurements can be converted to confidence intervals using the Neyman confidence belt, as described in Section 40 of the Particle Data Group handbook [12]. For Gaussians this is trivial. In frequentist usage, to say “I have measured $a$ to be $x\pm 3$.” means “I have a measurement $x$ from a device which returns values distributed about the true value $a$ according to a Gaussian distribution with a standard deviation of 3: I therefore assert that $x$ falls within the interval $[a-3,a+3]$ with probability 68%, which in the Gaussian case can be turned into saying that, given $x$, the interval $[x-3,x+3]$ covers the true value $a$ with 68% confidence. (It does not mean “The true value of $a$ lies within within $x-3$ and $x+3$ with 68% probability.”, that would be a Bayesian credible interval, though this subtle distinction may not always be appreciated.)
@@
If the pdf is not Gaussian – and for asymmetric errors this must be the case – then this trivial equivalence breaks. ‘Pdf errors’ describe the standard deviation and ‘likelihood errors’ the 68% central confidence interval. To take an artificially exaggerated but hopefully illuminating example, suppose a measuring device has a 50% chance of recording a measurement below the true value and does so according to a half-Gaussian with $\sigma=1$, and a 50% chance of recording a measurement above the true value, according to a half-Gaussian with $\sigma=5$. (This is the dimidiated Gaussian, described later.) If it measures $x=11.7$ this would be written as ${11.7}^{+5.0}_{-1.0}$, to describe the pdf. However the Neyman construction for the confidence belt would be ${11.7}^{+1.0}_{-5.0}$: if 11.7 is a downward fluctuation it cannot have come very far from the true value, whereas an upward fluctuation may have been produced by a true value ome distance away. This is shown in Figure 2.
@@
If we are considering a pdf $p(x;a)$ then we can integrate it to obtain the spread (as given by the variance) of the measured $x$ for a given value of $a$ and the ‘errors’ are expression of that spread. It tells us nothing about $a$, as that is given. If we are dealing with a likelihood function $L(a;x)$ we can get the ‘best’ estimate $\hat{a}$ and take the 68% CL central region from the $\Delta\ln L=-{1\over 2}$ points, [^1] This is only an approximation to the Neyman construction, but this point is generally not considered. For a discussion see Reference [13]. but we can say nothing more about $x$. It is true that a pdf can give 68% central limits, but in the language of the Neyman confidence belt they are the horizontal lines constructed on the plot it shows in Figure 2, not the vertical ones from which the confidence interval for a result is read off.So an ‘asymmetric error’ may refer to the variance and skewness of a pdf, or to the confidence region for a likelihood. Both cases arise in practice. We will therefore consider asymmetric pdfs and the variances (and higher moments) of $x$, and also, separately, asymmetric likelihoods and the confidence regions for $a$.
@@
One then has to consider how the ‘errors’ are going to be used. Again there are two possibilities: Combination of Errors and Combination of Results. This time the distinction is not so clear-cut.
@@ 1.2.1 Combination of Errors
This simple formula is where most physicists first meet statistics, in the context of experimental errors and uncertainties. For some function $u=u(x,y)$ where $x$ and $y$ are statistically independentwith errors that are small (so that a linear approximation is valid),
@@
where the two random variables $X_{1}$ and $X_{2}$ are supposed independent. If the pdfs are Gaussian, as shown in the left hand plot of Figure 3, where the dotted lines are the lines of constant $x_{2}-x_{1}$, then the integral gives a Gaussian for $\ell$. This Gaussian has standard deviation $\sqrt{\sigma_{1}^{2}+\sigma_{2}^{2}}$, as variances add. This is what is expressed by Equation (2).
@@
Combination of errors is usually a matter for pdfs, because variances add, even for non-Gaussian pdfs. This is dealt with in Section 2.2. A typical analysis will consider many sources of error, usually labelled as systematic, and the combination of these is a major concern of this paper. In such cases the Central Limit Theorem can be useful as when a large number of uncertainties is combined the overall distribution may be adequately described by a Gaussian, even though several of the contributions are asymmetric. However there are instances where combination of errors needs to be done using likelihoods and this is considered in Section 3.4.
@@ 1.2.2 Combination of Results
The second use we need to consider is the combination of results, also known as meta-analysis. Given a set of measurements of the same quantity one wishes to find the appropriate combined best value and its error – and usually one would also want some goodness of fit statistic to describe whether the different results are compatible. This can readily arise when the results are presented in the form of likelihoods (or the maximum likelihood estimator and the 68% central interval obtained from the likelihood function). It is conceptually simple to form the complete likelihood as the product of the individual ones, and find the location of the maximum and the 68% confidence band; there are technical challenges, and Section 3 deals with those. And this is the usual form of the problem. If these are all simple Gaussian measurements this requires the maximisation of $-{1\over 2}\sum_{i}\left({a_{i}-\hat{a}\over\sigma_{i}}\right)^{2}$, which occurs at $\hat{a}={\sum a_{i}/\sigma_{i}^{2}\over\sum 1/\sigma_{i}^{2}}$. (The index $i$ runs over the results being combined; each result may involve many data points but that is not relevant here.)
@@
But it is also possible that one has a set of results presented as pdfs: the $\hat{a}_{i}$ are estimators of the true $a$, they are functions of the data $x_{i}$ and are therefore described by pdfs. If they are unbiassed then any weighted sum $\sum_{i}w_{i}\hat{a}_{i}$ is an unbiased estimate, and one has the freedom to choose the $w_{i}$ to minimise the variance on this, using the combination of errors formula, Equation (2). For Gaussian errors this gives the well known prescription $w_{i}={1/\sigma_{i}^{2}\over\sum_{j}1/\sigma_{j}^{2}}$: the weight is proportional to the inverse square of the error. Hence for Gaussian errors the combination results using pdfs and using likelihoods gives the same answer, even though the meaning of the answer is subtly different. Again, for non-Gaussian errors this is not the case and we must consider both separately.
1.3 Summary on combination-of-error and combination-of-results
@@
The distinction between combination of errors and combination of results turns out to require clarification, perhaps because both are considered as ‘the error analysis’. The fundamental difference is that, in the first case, we want to quantify the precision of a combination, which involves different quantities, while in the second case, we compute the precision of an updated measurement of the same quantity, we obtain by combining different measurements of the same quantity. In the first case, precision goes down, because uncertainty adds. In the second case, precision goes up because we measure something with more data. (In statistics, the same name is kept for the precision, but different names are used for the outcome: measurement in the first case, estimate in the second.)
@@
Combination of errors fits naturally into the framework of pdf errors, though combination of results is perfectly possible where the result is a weighted sum of measured values. Conversely combination-of-results emerges very naturally from likelihoods, although combination of errors is perfectly possible through profiling.
For clarification the distinctions are shown in Table 1.
Combination of/Through PDF (Section 2) LIKELIHOOD (Section 3) Notes
@@ 2.1 Modelling pdf errors
The question as to how a model function, whether considered as a pdf or as a likelihood, should best be written in a non-Gaussian way is a very open one. In this section we consider pdf densities: likelihoods will be dealt with in Section 3. If we restrict ourselves to distributions which look similar to a Gaussian, we can expect them to require an additional third parameter, expressing the asymmetry in some way: we do not consider extensions to more than 3 parameters, but the formalism is open should they be needed.
@@
    the quantiles. The median $M$ is the middle quantile $x_{0.50}$, and the 68% (one sigma) central confidence probability region $[x_{0.16},x_{0.84}]$ can be written as $[M-{\sigma^{-}},M+{\sigma^{+}}]$. This can be expressed in the form ${M}^{+{\sigma^{+}}}_{-{\sigma^{-}}}$ and will be referred to as “the quantile parameters” in what follows. (The region is actually bounded by the values of the cumulative Gaussian $\Phi(-1)$ and $\Phi(+1)$: though here and elsewhere only 2 significant figures are shown the exact one sigma region is always meant.)
@@ 2.1.1 Flipped distributions
Figure 5: The top row shows an OPAT analysis where the deviations are both positive, though of differing magnitudeand the dimidiated and distorted fits. The second plot shows the dimidiated distribution: the total (black) is the sum of the two half-Gaussians (green and blue). The third plot also shows this pdf, together with the pdf obtained from the same 3 points using the distorted model, in green, and that from the equivalent pdf in red. The bottom row shows the equivalent for two negative deviations.
@@
This is implausible but not impossible. If the error concerned is important, then this presents a major problem: clearly something significant is going on, and the standard use of linear approximations in evaluating errors is inappropriate. Fortunately such cases are rare. A more typical situation where this arises, and it does arise, is in the evaluation of the contribution of many systematic uncertainties to the total error. Among the many small contributions there may be one or two where both deviations go the same way, though this is often due to statistical fluctuations on a small and unimportant quantity whose treatment will make no difference to the final quoted result. Nevertheless one needs a way of dealing gracefully with such cases. The best advice to the user is to repeat the analysis with more points and more data — but this extra effort may not be justifiable.
@@
The distorted model can handle such a situation. The minimum or maximum is now inside the 68% central confidence region so the need to consider both solutions in the pdf is very significant and the Jacobian peak is considerable, but all the algebra holds provided care is taken of the signs of ${\sigma^{-}}$ and ${\sigma^{+}}$.
@@
For the dimidiated model the distribution in $R$ is the sum of two half-Gaussians, with the same nominal mean, which is actually the extreme or cut-off value ${\cal M}$. If the ‘errors’ are $\sigma_{1}$ and $\sigma_{2}$ then, from the properties of the Gaussian, the combined distribution has moments
$\mu={\cal M}+{\sigma_{1}+\sigma_{2}\over\sqrt{2\pi}}\quad V={\sigma_{1}^{2}+\sigma_{2}^{2}\over 2}-{(\sigma_{1}+\sigma_{2})^{2}\over 2\pi}\quad\gamma=2{\sigma_{1}^{3}+\sigma_{2}^{3}\over\sqrt{2\pi}}-3{\sigma_{1}^{2}+\sigma_{2}^{2}\over 2}{\sigma_{1}+\sigma_{2}\over\sqrt{2\pi}}+2\left({\sigma_{2}+\sigma_{2}\over\sqrt{2\pi}}\right)^{3}.$ (4)
@@
It would be possible to introduce a specific flipped-dimidiated model using this algebra, but this seems an unnecessary complication to handle a situation which is either so serious that a simple 3-point analysis is inadequate, or so trivial that its impact will be small. We suggest instead that when a flipped distribution is encountered, it be replaced by a conventional dimidiated distribution with the same moments as given by Equation (4), and that the software (Appendices CD) should provide a tool for doing this. This, the red curve in Figure 5, has the same mean, variance and skewness as the black curve: they may appear different to the eye but this difference is not as large as it looks, and is similar to the differences between the dimidiated and distorted model (in green).
@@
Another scenario in which this formalism might be applied can occur with a discrete nuisance parameter; for example one might need to consider both the (favoured) normal and the (disfavoured but possible) inverted neutrino mass hierarchy. Suppose a result $R_{1}$ is obtained from the basic model or assumption, whereas a less-favoured alternative gives $R_{2}$. If, say $R_{2}>R_{1}$, one might wish to quote this as ${R_{1}}^{+R_{2}-R_{1}}_{-0}$. We are not recommending this, but if someone chooses to do so we can accommodate it as a flipped pdf where ${\sigma^{-}}$ and ${\sigma^{+}}$ are equal (or opposite: the signs require careful handling in the software) and given by $|R_{1}-R_{2}|$.

### v3

@@ 1.1 Background and motivation
Results in particle physics are often given with errors which are asymmetric, of the form ${R}^{+{\sigma^{+}}}_{-{\sigma^{-}}}$or instance, taking some results from the latest (at the time of writing) EPS conference:
@@
    from maximum likelihood estimation with errors given by the $\Delta\ln L=-{1\over 2}$ method, and the log likelihood is not a symmetric parabola. This is also shown in Figure 1. (The $\Delta\ln L=-{1\over 2}$ method is an approximation, but its use is widespread, essentially universal, and in what follows this is what is meant by ‘likelihood errors’.)
@@
The third case does not hold conceptual problems as one has, in principle, full knowledge of the underlying distribution function. The challenges come with the first and second cases, where one has to handle the quoted value and errors with no further information about how they arose. This prompts us to consider more closely some fundamental facts about the nature of ‘errors’ which are normally hidden from view by the convenient properties of the Gaussian function.
@@ 1.2 Beyond the Gaussian
The difference between rms spread ‘pdf errors’ and confidence interval ‘likelihood errors’ requires clarification. They both refer to uncertainties, the former from random probability and the latter from statistical inference. The Poisson process provides a simple illustration (though for a discrete probability distribution rather than a continuous pdf): the former describes the uncertainty of predicting results from a particular true mean, the latter the uncertainty of inferring the true value of the mean from a particular result.
@@
Pdf errors arise especially through measurements and their errors in experimental work. At high school and university the physicist is firmly taught never to quote an experimental result as a single number $x$, but always to give its error $\sigma$. That $\sigma$ is the standard deviation of the Gaussian distribution, as given in (1), from which this value is a sample. If, for better accuracy, the mean of $N$ measurements is quoted then it is accompanied by the standard error $\sigma/\sqrt{N}$ which is the standard deviation of the pdf of $\overline{x}$. The well-known formula for the combination of errors, given below as Equation 2, known to statisticians as the delta method, builds on this. It describes the combination of measurement errors: if $I$ is sampled from a Gaussian with standard deviation $\sigma_{I}$, and $V$ with $\sigma_{V}$, then the resistance measurement $R=V/I$ is sampled from a Gaussian for which (assuming the function can legitimately be linearised) $\sigma_{R}^{2}=\sigma_{V}^{2}/I^{2}+V^{2}\sigma_{I}^{2}/I^{4}$. The quoted result, $V/I\pm\sigma_{R}$, is saying the value has been drawn from a Gaussian distribution of unknown mean and standard deviation $\sigma_{R}$.
[…]
@@
If the pdf is not Gaussian — and for asymmetric errors this must be the case — then this trivial equivalence breaks. ‘Pdf errors’ describe the standard deviation and ‘likelihood errors’ the 68% central confidence interval. To take an artificially exaggerated but hopefully illuminating example, suppose a measuring device has a 50% chance of recording a measurement below the true value and does so according to a half-Gaussian with $\sigma=1$, and a 50% chance of recording a measurement above the true value, according to a half-Gaussian with $\sigma=5$. (This is the dimidiated Gaussian, described later.) If it measures $x=11.7$ this would be written as ${11.7}^{+5.0}_{-1.0}$, to describe the pdf. However the Neyman construction for the confidence belt would be ${11.7}^{+1.0}_{-5.0}$: if 11.7 is a downward fluctuation it cannot have come very far from the true value, whereas an upward fluctuation may have been produced by a true value one distance away. This is shown in Figure 2.
@@
If we are considering a pdf $p(x;a)$, then we can integrate it to obtain the spread (as given by the variance) of the measured $x$ for a given value of $a$ and the ‘errors’ are expression of that spread. It tells us nothing about $a$, as that is given. If we are dealing with a likelihood function $L(a;x)$ we can get the ‘best’ estimate $\hat{a}$ and take the 68% CL central region from the $\Delta\ln L=-{1\over 2}$ points, [^1] This is only an approximation to the Neyman construction, but this point is generally not considered. For a discussion see Reference [13]. but we can say nothing more about $x$. It is true that a pdf can give 68% central limits, but in the language of the Neyman confidence belt they are the horizontal lines constructed on the plot shown in Figure 2, not the vertical ones from which the confidence interval for a result is read off. So an ‘asymmetric error’ may refer to the variance and skewness of a pdf, or to the confidence region for a likelihood. Both cases arise in practice. We will therefore consider asymmetric pdfs and the variances (and higher moments) of $x$, and also, separately, asymmetric likelihoods and the confidence regions for $a$.
@@
One then has to consider how the ‘errors’ are going to be used. Again there are two possibilities: Combination of Errors and Combination of Results.
@@ 1.2.1 Combination of Errors
The simple formula for combination of errors is how most physicists first meet statistics, in the context of experimental errors and uncertainties. For some function $u=u(x,y)$ where the errors on $x$ and $y$ are statistically independent andsmall (so that a linear approximation is valid),
@@
where the two random variables $X_{1}$ and $X_{2}$ are assumed independent. If the pdfs are Gaussian, as shown in the left hand plot of Figure 3, where the dotted lines are the lines of constant $x_{2}-x_{1}$, then the integral gives a Gaussian for $\ell$. This Gaussian has standard deviation $\sqrt{\sigma_{1}^{2}+\sigma_{2}^{2}}$, as variances add. This is what is expressed by Equation (2).
@@
Combination of errors is usually a matter for pdfs, because variances add, even for non-Gaussian pdfs. This is dealt with in Section 2.2. A typical analysis will consider many sources of error, usually labelled as systematic, and the combination of these is a major concern of this paper. In such cases the Central Limit Theorem can be useful as when a large number of uncertainties is combined the overall distribution may be adequately described by a Gaussian, even though several of the contributions are asymmetric. However, there are instances where combination of errors needs to be done using likelihoods, and this is considered in Section 3.4.
@@ 1.2.2 Combination of Results
The second use we need to consider is the combination of results, also known as meta-analysis. Given a set of measurements of the same quantity one wishes to find the appropriate combined best value and its error — and usually one would also want some goodness of fit statistic to describe whether the different results are compatible. This can readily arise when the results are presented in the form of likelihoods (or the maximum likelihood estimator and the 68% central interval obtained from the likelihood function). It is conceptually simple to form the complete likelihood as the product of the individual ones, and find the location of the maximum and the 68% confidence band; there are technical challenges, and Section 3 deals with those. And this is the usual form of the problem. If these are all simple Gaussian measurements this requires the maximisation of $-{1\over 2}\sum_{i}\left({a_{i}-\hat{a}\over\sigma_{i}}\right)^{2}$, which occurs at $\hat{a}={\sum a_{i}/\sigma_{i}^{2}\over\sum 1/\sigma_{i}^{2}}$. (The index $i$ runs over the results being combined; each result may involve many data points but that is not relevant here.)
@@
But it is also possible that one has a set of results presented as pdfs: the $\hat{a}_{i}$ are estimators of the true $a$, they are functions of the data $x_{i}$ and are therefore described by pdfs. If they are unbiased then any weighted sum $\sum_{i}w_{i}\hat{a}_{i}$ is an unbiased estimate provided $\sum_{i}w_{i}=1$, and one has the freedom to choose the $w_{i}$ to minimise the variance on this, using the combination of errors formula, Equation (2). For Gaussian errors this gives the well known prescription $w_{i}={1/\sigma_{i}^{2}\over\sum_{j}1/\sigma_{j}^{2}}$: the weight is proportional to the inverse square of the error. Hence for Gaussian errors the combination results using pdfs and using likelihoods gives the same answer, even though the meaning of the answer is subtly different. Again, for non-Gaussian errors this is not the case and we must consider both separately.
1.3 Summary on combination-of-errors and combination-of-results
@@
The distinction between combination of errors and combination of results turns out to require clarification, perhaps because both are considered as ‘the error analysis’. The fundamental difference is that, in the first case, we want to quantify the precision of a combination, which involves different quantities, while in the second case, we compute the precision of an updated measurement of the same quantity, which we obtain by combining different measurements of the same quantity. In the first case, precision goes down, because uncertainty adds. In the second case, precision goes up because we measure something with more data. (In statistics, the same name is kept for the precision, but different names are used for the outcome: measurement in the first case, estimate in the second.)
@@
Combination-of-errors fits naturally into the framework of pdf errors, though combination of results is perfectly possible where the result is a weighted sum of measured values. Conversely combination-of-results emerges very naturally from likelihoods, although combination of errors is perfectly possible through profiling.
For clarification, the distinctions are shown in Table 1.
Combining [-2ex] [-2ex]Through PDF (Section 2) LIKELIHOOD (Section 3) Notes
@@ 2.1 Modelling pdf errors
The question as to how a model function, whether considered as a pdf or as a likelihood, should best be written in a non-Gaussian way is a very open one. In this section we consider pdfs: likelihoods will be dealt with in Section 3. If we restrict ourselves to distributions which look similar to a Gaussian, we can expect them to require an additional third parameter, expressing the asymmetry in some way: we do not consider extensions to more than 3 parameters, but the formalism is open should they be needed.
@@
    the quantiles. The median $M$ is the middle quantile $x_{0.50}$, and the 68% (one sigma) central confidence probability region $[x_{0.16},x_{0.84}]$ can be written as $[M-{\sigma^{-}},M+{\sigma^{+}}]$. This can be expressed in the form ${M}^{+{\sigma^{+}}}_{-{\sigma^{-}}}$ and will be referred to as “the quantile parameters” in what follows. (The region is actually bounded by the values of the cumulative Gaussian $\Phi(-1)$ and $\Phi(+1)$: though here and elsewhere only 2 significant figures are shown, the exact 1 sigma region is always meant.)
@@ 2.1.1 Flipped distributions
Figure 5: The top row shows an OPAT analysis where the deviations are both positive, though of differing magnitude, and the dimidiated and distorted fits. The second plot shows the dimidiated distribution: the total (black) is the sum of the two half-Gaussians (green and blue). The third plot also shows this pdf, together with the pdf obtained from the same 3 points using the distorted model, in green, and that from the equivalent dimidiatedpdf in red. The bottom row shows the samefor two negative deviations.
@@
This is implausible but not impossible. If the error concerned is important then this presents a major problem: clearly something significant is going on, and the standard use of linear approximations in evaluating errors is inappropriate. Fortunately such cases are rare. A more typical situation where this arises, and it does arise, is in the evaluation of the contribution of many systematic uncertainties to the total error. Among the many small contributions there may be one or two where both deviations go the same way, though this is often due to statistical fluctuations on a small and unimportant quantity whose treatment will make no difference to the final quoted result. Nevertheless one needs a way of dealing gracefully with such cases. The best advice to the user is to repeat the analysis with more points and more data — but this extra effort may not be justifiable.
@@
The distorted model can handle such a situation. The minimum or maximum is now inside the 68% central confidence region so the need to consider both branches of the parabola forthe pdf is very significant and the Jacobian spike is considerable, but all the algebra holds provided care is taken of the signs of ${\sigma^{-}}$ and ${\sigma^{+}}$.
@@
For the dimidiated model the distribution in $R$ is the sum of two half-Gaussians, with the same nominal mean, which is actually the extreme or cut-off value ${\cal M}$. If the ‘errors’ are $\sigma_{1}$ and $\sigma_{2}$ and are both positive, as in the upper row of Figure 5,then, from the properties of the Gaussian, the combined distribution has moments
$\mu={\cal M}+{\sigma_{1}+\sigma_{2}\over\sqrt{2\pi}},\quad V={\sigma_{1}^{2}+\sigma_{2}^{2}\over 2}-{(\sigma_{1}+\sigma_{2})^{2}\over 2\pi},\quad\gamma=2{\sigma_{1}^{3}+\sigma_{2}^{3}\over\sqrt{2\pi}}-3{\sigma_{1}^{2}+\sigma_{2}^{2}\over 2}{\sigma_{1}+\sigma_{2}\over\sqrt{2\pi}}+2\left({\sigma_{2}+\sigma_{2}\over\sqrt{2\pi}}\right)^{3}.$ (4)
@@
It would be possible to introduce a specific flipped-dimidiated model using this algebra, but this seems an unnecessary complication to handle a situation which is either so serious that a simple 3-point analysis is inadequate, or so trivial that its impact will be small. We suggest instead that, when a flipped distribution is encountered, it be replaced by a conventional dimidiated distribution with the same moments as given by Equation (4), and that the software (Appendices CD) should provide a tool for doing this. This, the red curve in Figure 5, has the same mean, variance and skewness as the black curve: they may appear different to the eye but this difference is not as large as it looks, and is similar to the differences between the dimidiated and distorted model (in green).
@@
Another scenario in which this formalism might be applied can occur with a discrete nuisance parameter; for example one might need to consider both the (favoured) normal and the (disfavoured but possible) inverted neutrino mass hierarchy. Suppose a result $R_{1}$ is obtained from the basic model or assumption, whereas a less-favoured alternative gives $R_{2}$. If, say $R_{2}>R_{1}$, one might wish to quote this as ${R_{1}}^{+|R_{2}-R_{1}|}_{-0}$. We are not recommending this, but if someone chooses to do so we can accommodate it as a flipped pdf where ${\sigma^{-}}$ and ${\sigma^{+}}$ are equal (or opposite: the signs require careful handling in the software) and given by $|R_{1}-R_{2}|$.